Mary Julie and Peter went to restaurant for dinner. Mary paid 1/3 of the total cost. Julie paid Half as much as Mary did and Peter paid the remaining $75. What was the total cost of dinner?
Answer by mananth(16949)

let the cost of dinner be x
Mary paid x/3
Julie 1/2 * x/3= x/6
Balance $75 Peter paid
x/3 + x/6 +75 = x
x/3+x/6 -x= -75
LCD = 6
multiply equation by 6
2x+x-6x=-450
-3x=-450
/-3
x=150
Mary paid 1/3 of 150 = $50
Julie paid 1/6 of 150 = $25
Peter paid $75
Add up
The total cost of dinner = $150